Translate the following situation into an equation. Do not solve.
“The length of the second side of a triangle is four less than three times the length of the first side. The length of the third side is one more than the length of the first side. If the perimeter of the triangle is 37 feet, what is the length of the first side?”

Answer by stanbon(75887)

Do not solve.
“The length of the second side of a triangle is four less than three times the length of the first side. The length of the third side is one more than the length of the first side. If the perimeter of the triangle is 37 feet, what is the length of the first side?”
---
Equations:
Let the sides be x,y,z
---
y = 3x-4
z = x+1
x+y+z = 37
---
x + (3x-4)+(x+1) = 37
5x-3 = 37
5x = 40
x = 8
============
Cheers,
Stan H.
============